# Enhancing Clinical Trial Patient Matching through Knowledge Augmentation and Reasoning with Multi-Agents

### v2

@@ A.4. MAKAR-Augemented N2C2 Criteria
ABDOMINAL Criteria: History of intra-abdominal surgery, including small/large intestine resection or small bowel obstruction.
@@
ADVANCED-CAD Criteria: Advanced cardiovascular disease (CAD) defined by having two or more of the following: (a) Taking 2 or more medications for CAD, (b) History of myocardial infarction (MI), (c) Current angina, or (d) Ischemia (past or present).
@@
ALCOHOL-ABUSE Criteria: Current alcohol use exceeding weekly recommended limits.
Explanation: Weekly limits: up to 14 standard drinks. A standard drink is  14g of pure alcohol (12oz beer, 5oz wine, or 1.5oz spirits).
ASP-FOR-MI Criteria: Use of aspirin to prevent myocardial infarction (MI).
@@
CREATININE Criteria: Serum creatinine level above the upper normal limit.
Explanation: Creatinine reflects kidney function. Normal levels vary by age and gender (e.g.,  0.6-1.2 mg/dL for adult males). Elevated levels suggest potential kidney impairment.
DIETSUPP-2MOS Criteria: Consumption of a dietary supplement (excluding Vitamin D) in the past 2 months.
@@
DRUG-ABUSE Criteria: Current or past drug abuse history.
@@
ENGLISH Criteria: Patient must speak English.
@@
HBA1C Criteria: Hemoglobin A1c (HbA1c) between 6.5% and 9.5%.
@@
KETO-1YR Criteria: Diagnosis of ketoacidosis in the past year.
@@
MAJOR-DIABETES Criteria: Major diabetes-related complications like amputation, kidney damage, or neuropathy.
@@
MAKES-DECISIONS Criteria: Patient must make their own medical decisions.
@@
MI-6MOS Criteria: Myocardial infarction (MI) within the past 6 months.
@@
A.5. Sampled ClinicalTrial.gov Criteria
Inclusion Criteria Exclusion Criteria
1. Locally advanced or metastatic breast cancer confirmed by histopathology. 1. Patients known or suspected intolerance or hypersensitivity to main ingredient or any of the excipients of SNB-101.
@@
2. Received 2 previous lines of anti-tumor treatment and developed resistance to standard treatment. 2. Patients with bilateral invasive breast cancers.
@@
3. Patients with locally advanced or metastatic advanced solid tumors confirmed by histology or cytology (including but not limited to triple-negative breast cancer, gastric cancer, colorectal cancer); Patients with 1 line of standard treatment failure (disease progression after treatment or intolerability of toxic side effects of treatment), or no standard treatment, or unable to receive standard treatment. 3. Age $<$18 years old or $>$70 years old.
@@
4. At least 1 measurable lesion per Response Evaluation Criteria in Solid Tumors version 1.1 and previously treated lesions with radiotherapy or focal therapy and no progression cannot be included as target lesion for assessment. 4. Patients with standard metallic contraindications to CMR or estimated glomerular filtration rate $<$ 30 mL/min/1.73 m².
@@
5. Histologically confirmed breast cancer with an invasive component measuring 20 mm and/or with morphologically confirmed spread to regional lymph nodes (stage cT2-cT4 with any cN, or cN1-cN3 with any cT). 5. History of hypersensitivity or contraindication to TTF.
@@
6. Known HER2-positive breast cancer defined as an IHC status of 3+. If IHC is 2+, a positive in situ hybridization (FISH, CISH, or SISH) test is required by local laboratory testing. 6. Implanted pacemaker, defibrillator, or other electrical medical devices.
@@
7. Patients with breast cancer planned to receive Anthracycline and Cyclophosphamide chemotherapy. 7. Bilateral breast cancer (including multifocal breast cancer. All histological lesions were HER2 1+ or 2+, not detected by FISH amplification and HR positive according to ASCO guidelines).
@@
8. Patients treated for hormone-dependent localized breast cancer requiring adjuvant hormonal therapy (HT). 8. Received any form of anti-tumor therapy (chemotherapy, radiotherapy, molecular targeted therapy, endocrine therapy, etc.).
9. Patients treated with tamoxifen for a maximum of 1 to 3 years. 9. Breast cancer without histopathological diagnosis.
@@
10. Histologically confirmed ER+/HER2- early-stage resected invasive breast cancer at high or intermediate risk of recurrence, , based on clinical-pathological risk features, as defined in the protocol. 10. Known personal history of ductal carcinoma in situ (DCIS) or invasive breast cancer.
@@
11. Completed adequate (definitive) locoregional therapy (surgery with or without radiotherapy) for the primary breast tumour(s), with or without (neo)adjuvant chemotherapy. 11. Prior systemic treatment for any malignancy.
@@
12. Completed at least 2 years but no more than 5 years (+3 months) of adjuvant ET (+/- CDK4/6 inhibitor). 12. Significant medical comorbidities as per investigator evaluation.
13. Patients receiving letrozole for more than two months. 13. Patients with ongoing indications for the cardioprotective medication - ACE inhibitors, ARBs and/or beta-blockers
14. Postmenopausal hormone receptor-positive patients. 14. Intermediate or high-grade ductal carcinoma in situ.
@@
15. Patients with N0 disease or those who have undergone targeted axillary detection with a good response. 15. Invasive carcinoma.
Table 9. Sampled Criteria from Breast Cancer Clinical Trials

### v3

@@ B.3 N2C2 Criteria Refined by GPT-4o
Criteria: History of intra-abdominal surgery, including small/large intestine resection or small bowel obstruction.
@@
ADVANCED-CAD
Criteria: Advanced cardiovascular disease (CAD) defined by having two or more of the following: (a) Taking 2 or more medications for CAD, (b) History of myocardial infarction (MI), (c) Current angina, or (d) Ischemia (past or present).
@@
ALCOHOL-ABUSE
Criteria: Current alcohol use exceeding weekly recommended limits.
Explanation: Weekly limits: up to 14 standard drinks. A standard drink is approximately 14g of pure alcohol (12oz beer, 5oz wine, or 1.5oz spirits).
ASP-FOR-MI
Criteria: Use of aspirin to prevent myocardial infarction (MI).
@@
CREATININE
Criteria: Serum creatinine level above the upper normal limit.
Explanation: Creatinine reflects kidney function. Normal levels vary by age and gender (e.g., approximately 0.6–1.2 mg/dL for adult males). Elevated levels suggest potential kidney impairment.
DIETSUPP-2MOS
Criteria: Consumption of a dietary supplement (excluding Vitamin D) in the past 2 months.
@@
DRUG-ABUSE
Criteria: Current or past drug abuse history.
@@
ENGLISH
Criteria: Patient must speak English.
@@
HBA1C
Criteria: Hemoglobin A1c (HbA1c) between 6.5% and 9.5%.
@@
KETO-1YR
Criteria: Diagnosis of ketoacidosis in the past year.
@@
MAJOR-DIABETES
Criteria: Major diabetes-related complications like amputation, kidney damage, or neuropathy.
@@
MAKES-DECISIONS
Criteria: Patient must make their own medical decisions.
@@
MI-6MOS
Criteria: Myocardial infarction (MI) within the past 6 months.
@@
B.4 Criteria from ClinicalTrials.gov
Inclusion Criteria
1. Locally advanced or metastatic breast cancer confirmed by histopathology.
@@
2. Received 2 previous lines of anti-tumor treatment and developed resistance to standard treatment.
@@
3. Patients with locally advanced or metastatic advanced solid tumors confirmed by histology or cytology (including but not limited to triple-negative breast cancer, gastric cancer, colorectal cancer); Patients with 1 line of standard treatment failure (disease progression after treatment or intolerability of toxic side effects of treatment), or no standard treatment, or unable to receive standard treatment.
@@
4. At least 1 measurable lesion per Response Evaluation Criteria in Solid Tumors version 1.1 and previously treated lesions with radiotherapy or focal therapy and no progression cannot be included as target lesion for assessment.
@@
5. Histologically confirmed breast cancer with an invasive component measuring 20 mm and/or with morphologically confirmed spread to regional lymph nodes (stage cT2-cT4 with any cN, or cN1-cN3 with any cT).
@@
6. Known HER2-positive breast cancer defined as an IHC status of 3+. If IHC is 2+, a positive in situ hybridization (FISH, CISH, or SISH) test is required by local laboratory testing.
@@
7. Patients with breast cancer planned to receive Anthracycline and Cyclophosphamide chemotherapy.
@@
8. Patients treated for hormone-dependent localized breast cancer requiring adjuvant hormonal therapy (HT).
9. Patients treated with tamoxifen for a maximum of 1 to 3 years.
@@
10. Histologically confirmed ER+/HER2- early-stage resected invasive breast cancer at high or intermediate risk of recurrence, based on clinical-pathological risk features, as defined in the protocol.
@@
11. Completed adequate (definitive) locoregional therapy (surgery with or without radiotherapy) for the primary breast tumour(s), with or without (neo)adjuvant chemotherapy.
@@
12. Completed at least 2 years but no more than 5 years (+3 months) of adjuvant ET (+/- CDK4/6 inhibitor).
13. Patients receiving letrozole for more than two months.
14. Postmenopausal hormone receptor-positive patients.
@@
15. Patients with N0 disease or those who have undergone targeted axillary detection with a good response.
Exclusion Criteria
1. Patients known or suspected intolerance or hypersensitivity to main ingredient or any of the excipients of SNB-101.
2. Patients with bilateral invasive breast cancers.
3. Age $<$18 years old or $>$70 years old.
4. Patients with standard metallic contraindications to CMR or estimated glomerular filtration rate $<$30 mL/min/1.73 m².
5. History of hypersensitivity or contraindication to TTF.
6. Implanted pacemaker, defibrillator, or other electrical medical devices.
7. Bilateral breast cancer (including multifocal breast cancer. All histological lesions were HER2 1+ or 2+, not detected by FISH amplification and HR positive according to ASCO guidelines).
8. Received any form of anti-tumor therapy (chemotherapy, radiotherapy, molecular targeted therapy, endocrine therapy, etc.).
9. Breast cancer without histopathological diagnosis.
10. Known personal history of ductal carcinoma in situ (DCIS) or invasive breast cancer.
11. Prior systemic treatment for any malignancy.
12. Significant medical comorbidities as per investigator evaluation.
13. Patients with ongoing indications for the cardioprotective medication - ACE inhibitors, ARBs and/or beta-blockers.
14. Intermediate or high-grade ductal carcinoma in situ.
15. Invasive carcinoma.
B.5 Example Criteria from ClinicalTrials.gov refined by GPT-4o
Criterion: Locally advanced or metastatic breast cancer confirmed by histopathology.
Additional Explanation:
- Locally advanced breast cancer refers to cancer that has spread beyond the breast to nearby areas but not to distant body parts. It often involves large tumors or extensive lymph node involvement.
- Metastatic breast cancer means that the cancer has spread from the breast to other parts of the body, such as the bones, liver, lungs, or brain.
- Histopathology is the examination of tissues under a microscope to study the manifestations of disease. This confirms the diagnosis of breast cancer by analyzing tissue samples obtained from a biopsy or surgery.
Criterion: Received 2 previous lines of anti-tumor treatment and developed resistance to standard treatment.
Additional Explanation:
1. **Anti-tumor treatment**: This refers to any therapy aimed at treating cancer, which might include chemotherapy, targeted therapy, immunotherapy, or hormone therapy.
2. **2 previous lines of treatment**: In cancer therapy, a "line" of treatment refers to a complete sequence of therapy given for cancer, which can consist of one or multiple medications or interventions. The patient must have undergone two different complete treatment sequences.
3. **Developed resistance**: This means that the cancer no longer responds to the standard treatment. This can be defined by the tumor growing or not shrinking in size despite treatment, or the cancer markers take not being reduced.
4. **Standard treatment**: This refers to the most widely accepted and utilized treatments for a particular type of cancer, as established by clinical research and guidelines. Examples might include first-line chemotherapy regimens or targeted therapies. Exact "standard treatment" could vary based on cancer type, but usually involves those strategies that have the highest success rates based on evidence.
Appendix C Prompts
C.1 The prompt for zero-shot matching
You are an experienced physician tasked with determining whether a patient meets specific eligibility criteria for clinical trials. Below, I will provide multiple EHR records for a single patient along with one criterion. You must evaluate whether the patient satisfies the criterion. For each criterion, respond strictly with 1 if the criterion is met, or 0 if it is not met. If there is not enough information to make a determination, return 0. Respond with only 0 or 1; no reasoning or explanation is needed.
Patient EHR Data:…
The Ctiterion:…
C.2 The prompt for refining criteria
You are an experienced physician tasked with reviewing the trial criterion to make it easy to follow for less-trained staff. Sometimes, the criterion may seem vague. If it is, please complement the criterion for clarity in plain English by: (1) defining any specialized medical terms, and (2) explaining vague phrases such as ‘normal levels’ by specifying normal ranges for different genders and age groups. You must keep the original criterion intact and add additional explanations only if you think they are necessary. Please ONLY return the revised criterion and use the following format:
<Original Criterion>: XXX
<Additional Explanation>: XXX
Here is the original criterion:…
C.3 Prompts for Agents
C.3.1 Router Agent
You are a routing agent responsible for selecting the most appropriate augmentation agent to enrich a trial criterion with relevant, domain-specific information. You will receive one criterion as input. Based on its content, determine which augmentation agent to use.
**Retrieval Agent**: Use this agent if the criterion requires authoritative information from indexed medical databases like MedDRA and FDALabel.
**Online Search Agent**: Use this agent if the criterion needs supplementary knowledge from external online sources.
**Self-Augment Agent**: Use this agent if the criterion can be refined using the language model’s intrinsic medical knowledge.
Strictly respond with only one of the following labels indicating your choice: Retrieval, Online Search, or Self-Augment. No reasoning or explanation is needed—only return the label.
The Criterion:..
C.3.2 Knowledge Augmentation Agent
You are an experienced physician tasked with reviewing the trial criterion to make it easy to follow for less-trained staff. Sometimes, the criterion may seem vague. If it is, please complement the criterion for clarity in plain English by: (1) defining any specialized medical terms, and (2) explaining vague phrases such as ‘normal levels’ by specifying normal ranges for different genders and age groups. You must keep the original criterion intact and add additional explanations only if you think they are necessary. Please ONLY return the revised criterion and use the following format:
<Original Criterion>: XXX
<Additional Explanation>: XXX
Here is the original criterion:…
C.3.3 Critic Agent
You are a senior physician reviewing criteria revised by another physician. Make sure that (1) the original criterion is preserved, and (2) any additional explanations are accurate. For each criterion, respond strictly with **1** if both checks pass, or **0** if they do not. No reasoning or explanation is needed.
The Revised Ctiterion:…
The Original Ctiterion:…
C.3.4 Reasoning Agent
You are a Principal Investigator (PI) in a clinical trial. You have multiple EHR records for one patient and a criterion with detailed explanation. Please perform a reasoning process to analyze why this patient meets or does not meet this trial criterion. If the information is not sufficient to make a determination, assume the patient does not meet the criterion.
Patient EHR Data:…
The Detailed Ctiterion:…
C.3.5 Matching Agent
You are a Principal Investigator (PI) in a clinical trial and are tasked with determining whether a patient meets specific eligibility criteria. Below, I will provide multiple EHR records for one patient, one criterion, and a reasoning process from another PI. You must evaluate whether the patient meets the criterion. For each criterion, respond strictly with **1** if met or **0** if not met. Respond with 0 or 1 only; no reasoning or explanation is needed.
Patient EHR Data:…
The Ctiterion:…
The Reasoning Process:…
Appendix D Pseudo-Code for MAKAR Workflow
Algorithm 1 MAKAR: Multi-Agent Knowledge Augmentation and Reasoning
1:Original criterion $C_{orig}$, Patient EHR data $P_{EHR}$
2:Matching result $M\in\{\text{0},\text{1}\}$
3:// Augmentation Module: MAKAR-Aug
4:$C_{aug}\leftarrow\text{RouterAgent}(C_{orig})$
5:if RouterAgent determines $\text{RouterAgent determines }C_{orig}\text{ is detailed enough}$ then
6:     $C_{aug}\leftarrow C_{orig}$triangleright Forward directly to Reasoning Module
7:else
[…]
10:         $C_{aug}\leftarrow\text{RetrievalAgent}(C_{orig},\text{databases})$
[…]
12:         $C_{aug}\leftarrow\text{OnlineSearchAgent}(C_{orig})$
[…]
14:         $C_{aug}\leftarrow\text{SelfAugmentAgent}(C_{orig})$
15:     end if
16:     // Iterative refinement with Critic Agent
17:     repeat
18:         $\text{feedback}\leftarrow\text{CriticAgent}(C_{aug},C_{orig})$
19:         if feedback.rejected then
20:              $C_{aug}\leftarrow\text{Re-augment}(C_{orig},\text{feedback})$
21:         end if
22:     until CriticAgent accepts $\text{CriticAgent accepts }C_{aug}$
23:end if
24:// Reasoning Module: MAKAR-Reason
[…]
26:// Generate detailed evaluation
27:for each concept $c$ in $C_{aug}$ do
28:     $\text{evaluation}[c]\leftarrow\text{ReasoningAgent.evaluate}(c,P_{EHR})$
29:end for
30:// Final matching decision
[…]
32:// Supervision validation
[…]
34:if validation.approved then
35:     return $M$
36:else
37:     return validation.corrected_result
38:end if